OpenStudy (anonymous):

Someone please help me. i wasted an hour trying to get 2 other questions and i am now behind. i need really quick answers now. so please just put the answer and then the explaination. Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3) Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8 Choose

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (shubhamsrg):

can you show me your work ? ;)

OpenStudy (anonymous):

uhm excuse me ?

OpenStudy (shubhamsrg):

nothing,,nevermind..

OpenStudy (anonymous):

what ? i meant no attitude by that, i dont know what u mean ?

OpenStudy (anonymous):

What's so confusing about these questions?

OpenStudy (anonymous):

i dont know how to do them and i am in a real big hurry.

OpenStudy (anonymous):

So, for the first one, rewrite the equation in slope intercept form. You have x + 7y = -21

OpenStudy (cwrw238):

x + 7y = -21 x-intercept : put y = 0 and calculate x + 0 = -21 y-intercept: put x = 0 and work out y

OpenStudy (cwrw238):

the x intercept is -21

OpenStudy (anonymous):

@zordoloom i dont see that in the answer choices ????

OpenStudy (cwrw238):

* (-21, 0)

OpenStudy (anonymous):

When you rewrite it , you get the y intercept at (0, -3) and the x at (-21,0)

OpenStudy (cwrw238):

y-intercept : when x = 0 is x0 + 7y = -21 y = -3 thats (0,-3)

OpenStudy (anonymous):

okay i got the first one. how bout the next one

OpenStudy (anonymous):

and thank you both @cwrw238 @zordoloom for helping me !!

OpenStudy (anonymous):

Use point slope form. Use the slope from the first equation, and plug in the point that is given (6, -2) into the equation

OpenStudy (anonymous):

since the slope is just 1, it shouldn't be that complicated to write the equation.

OpenStudy (cwrw238):

plug in x=6 and y=-2 into each equation and see if they fit eg try y = x + 4 -2 = 6 + 4 -2 = 10!!!! - so its not that one

OpenStudy (cwrw238):

try the last one

OpenStudy (anonymous):

y-(-2)=1(x-6), just rearrange it.

OpenStudy (anonymous):

y=x-8

OpenStudy (anonymous):

And for the last one, just use the inverse slope of the given equation to find the line that is perpendicular.

OpenStudy (anonymous):

i think the last one is y = x + 2

OpenStudy (anonymous):

dat right ?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

idk who to give da best medal to ?

OpenStudy (cwrw238):

no - the last one is y = x - 8 - what zordoloom got.

OpenStudy (cwrw238):

i gave it to to zordollom already

OpenStudy (cwrw238):

oh - ok - i'll take a medal as well l lol!!

OpenStudy (anonymous):

So for the third one, (y-(-2)=1(x-(-4))

OpenStudy (anonymous):

So, ya the third one is correct y=x+2

OpenStudy (anonymous):

okay haha thank u both sooooooo much !! u have no idea how much u helped :)